OpenStudy (anonymous):

At a dairy it takes 90 seconds to fill 30 one-gallon jugs of milk. Howmany minutes does it take to fill 90 jugs of milk

OpenStudy (anonymous):

90 jugs of milk will take 3 times longer i.e. 3*90/60 = 270/60 = 27/6 = 4.5 miuntes

OpenStudy (texaschic101):

90 seconds = 1.5 minutes 1.5 / 30 = x / 90 --(1.5 min to 30 jugs = x min to 90 jugs) cross multiply 30x = 135 x = 135/30 x = 4.5 minutes